(1)To balance a 37.0-kg automobile tire and wheel, a mechanic must place a 55.0-g lead weight 21.0cmfrom the center of the wheel. When the wheel is balanced, its center of mass is exactly at the center of the wheel.

How far from the center of the wheel was its center of mass before the lead weight was added?

(2)

How far from the center of the wheel was its center of mass before the lead weight was added?

(A)

If the car was in neutral and its brakes were off, so that the collision is approximately elastic, find the final speed of the truck.

(B)

Find the final speed of the car.
Answer #1

when the beam is balanced, Moments are conserved

so

m ds = MS

so

0.055 * 0.21 = 37 * x

x= 0.055 *0.21/37

x = 0.0312 cm

------------------------------

data seems missing for part B and C